# Exploring Network-Knowledge Graph Duality: A Case Study in Agentic Supply Chain Risk Analysis

### v1

@@ 1. Introduction
Large language models (LLMs) are increasingly embedded in critical, high-stakes decision processes, from medical diagnostic triage (Dedhia et al., 2025) to financial advice(Takayanagi et al., 2025). However, most domain deployments still rely on fine‑tuning a specialist model on a hand‑curated corpus or on straightforward retrieval augmented generation (RAG) that proxies relationships with vector distance. Both paths leave value on the table for supply chain risk analysis: the former is costly to update with current events and freezes knowledge at training time, while the latter ignores the rich network semantics hidden in structured data. In addition, they typically operate on a single modality, despite the fact that domain‑specific tasks require seamless reasoning over text, tables, graphs, and time‑series data.
@@ 2. Background and Related Work
Recent work on bottom‑up domain‑specific super‑intelligence (BDSI) shows how multi‑hop KG paths can be verbalized into 24,000 reasoning tasks that supervise a 32 B‑parameter model, yielding state‑of‑the‑art scores on ICD‑Bench (Dedhia et al., 2025). Their curriculum transforms the graph structure into training data, pushing the model to compose primitives into higher‑order concepts.
@@
GraphRAG (Edge et al., 2024) first extracts an entity knowledge graph from a document corpus, then offline generates hierarchical community summaries using a Leiden‑based clustering; at query time, it assembles partial answers from those summaries and fuses them into a global response. This excels on corpus‑wide “sense‑making” questions but its pre‑processing and multi‑stage summarization pipeline are compute‑intensive. Lighter variants such as Neural‑KB, GNN‑RAG(Mavromatis and Karypis, 2024) and temporal‑aware RAG(Zhu et al., 2025) likewise keep the base LM frozen while injecting graph structure to boost recall and freshness. For real-time applications such as finance, retraining such a model daily is even less feasible.
@@
Platforms such as Neo4j,(Neo4j Inc., 2025) and LangChain’s Graph Retriever,(Bratanic, 2024) execute a pattern like
@@
Our proposed agentic framework bridges the gap: it agrees with the insight in (Dedhia et al., 2025) that KG paths encode domain reasoning, but applies that insight at inference time with lightweight network traversal, avoiding the cost and staleness of specialist fine‑tuning while remaining faster than on‑the‑fly graph construction methods like GraphRAG.
@@ 3. System Overview
Upon initiating a chat, the user supplies our system with the portfolio constituents and their weights. We make sure that the LLM holds this at the top of the context window to always be able to refer to it, since LLMs are known to get “lost in the middle” and not pay attention to data in the middle of the context window(Liu et al., 2023).
@@
The selected tools tap separate FAISS indices(Douze et al., 2024), one per modality, so semantic relationships inside each data type remain intact and compound queries [^1] A request is termed a compound query when it references two or more distinct sub-topics or entities. Embedding the full sentence typically positions its vector between the relevant clusters in space, so a $k$-nearest search can miss evidence dispersed across those clusters. For example, in the query “factor exposures and cobalt” there may be no single source covering both topics together, while each topic is well covered individually at a significant vector distance from the other; $k$-nearest retrieval therefore fails to return comprehensive coverage. can be resolved cleanly across stores. Each tool call is executed through the OpenAI function-calling interface; the language model emits a JSON stub, the back-end runs the call, and the retrieved snippets flow into a transient database before being re-injected into the prompt. Finally, a frozen GPT-4o model synthesizes those multi-modal fragments with the original question, streams an Augmented Response to the UI, and stores the turn in memory, readying the loop for whatever the user asks next. Responses are fast, as the tool calling system is a series of quick API calls.
@@ 4. Data Modes And Tools
For every security in the user’s portfolio, we ingest MSCI MAC(Shepard et al., 2020) factor scores and the accompanying methodology text. Before embedding, each record is wrapped in a context shell (see section 5): a short paragraph that “tokenizes” the numerical z‑scores within natural‑language sentences so the LLM can reason about their meaning rather than seeing them as opaque numbers. The resulting embeddings power the get_factors tool.
@@
The final mode is a multi‑entity knowledge graph that links Company, Product, Input Product, Input to Input Product, Industry, and Location nodes via relations such as Produces, Has Input, Manufactured In and so forth (the full list of edge types are given in LABEL:tab:entity_rships_kg in A. Although illustrative rather than fully industrial, the graph is constructed with a pipeline akin to AIPNet (Fetzer et al., 2024), to query ChatGPT’s intrinsic knowledge base to create a dataset of the desired form. These paths enter the prompt as restructured text within a context shell, exposing the semantic relationships between nodes on the traversed graph.
@@ 5. Context Shell
Domain risk analysis is based on numbers: z-scored factor exposures, portfolio weights, valuation ratios. Large language models, however, treat raw figures as opaque tokens and rarely attach contextual economic meaning to them. To retain a frozen, generic LLM while still letting it reason quantitatively, our proposed system wraps every table row from the MSCI Multi-Asset-Class (MAC) factor model in a context shell: a short paragraph that “tokenizes” each figure inside an explanatory sentence so that neighboring words endow the number with semantics. The surrounding text becomes part of the embedding; the number itself is now a first-class token that via attention relationships with pertinent context becomes easily interpretable. A related but different idea has been explored in (Anthropic, 2024).
@@ 6. Network‑Science Path Discovery
In practice supply chain networks follow a scale-free, power-law degree distribution (Mungo et al., 2023): hubs such as Electronic Integrated Circuits in the middle of Figure 3 in the Apple network expose a wealth of information in just one hop, whereas peripheral nodes require additional traversal hops for useful coverage.

### v2

@@ 1. Introduction
Large language models (LLMs) are increasingly embedded in critical, high-stakes decision processes, from medical diagnostic triage (dedhia2025bupsi) to financial advice(takayanagi2025generative). However, most domain deployments still rely on fine‑tuning a specialist model on a hand‑curated corpus or on straightforward retrieval augmented generation (RAG) that proxies relationships with vector distance. Both paths leave value on the table for supply chain risk analysis: the former is costly to update with current events and freezes knowledge at training time, while the latter ignores the rich network semantics hidden in structured data. In addition, they typically operate on a single modality, despite the fact that domain‑specific tasks require seamless reasoning over text, tables, graphs, and time‑series data.
@@ 2. Background and Related Work
Recent work on bottom‑up domain‑specific super‑intelligence (BDSI) shows how multi‑hop KG paths can be verbalized into 24,000 reasoning tasks that supervise a 32 B‑parameter model, yielding state‑of‑the‑art scores on ICD‑Bench (dedhia2025bupsi). Their curriculum transforms the graph structure into training data, pushing the model to compose primitives into higher‑order concepts.
@@
GraphRAG (edge2024graphrag) first extracts an entity knowledge graph from a document corpus, then offline generates hierarchical community summaries using a Leiden‑based clustering; at query time, it assembles partial answers from those summaries and fuses them into a global response. This excels on corpus‑wide “sense‑making” questions but its pre‑processing and multi‑stage summarization pipeline are compute‑intensive. Lighter variants such as Neural‑KB, GNN‑RAG(mavromatis2024gnn) and temporal‑aware RAG(zhu2025fintmmbench) likewise keep the base LM frozen while injecting graph structure to boost recall and freshness. For real-time applications such as finance, retraining such a model daily is even less feasible.
@@
Platforms such as Neo4j,(neo4j_cypher_manual) and LangChain’s Graph Retriever,(bratanic2024langchain) execute a pattern like
@@
Our proposed agentic framework bridges the gap: it agrees with the insight in (dedhia2025bupsi) that KG paths encode domain reasoning, but applies that insight at inference time with lightweight network traversal, avoiding the cost and staleness of specialist fine‑tuning while remaining faster than on‑the‑fly graph construction methods like GraphRAG.
@@ 3. System Overview
Upon initiating a chat, the user supplies our system with the portfolio constituents and their weights. We make sure that the LLM holds this at the top of the context window to always be able to refer to it, since LLMs are known to get “lost in the middle” and not pay attention to data in the middle of the context window(liu2023lostmiddlelanguagemodels).
@@
The selected tools tap separate FAISS indices(douze2024faiss), one per modality, so semantic relationships inside each data type remain intact and compound queries [^1] A request is termed a compound query when it references two or more distinct sub-topics or entities. Embedding the full sentence typically positions its vector between the relevant clusters in space, so a $k$-nearest search can miss evidence dispersed across those clusters. For example, in the query “factor exposures and cobalt” there may be no single source covering both topics together, while each topic is well covered individually at a significant vector distance from the other; $k$-nearest retrieval therefore fails to return comprehensive coverage. can be resolved cleanly across stores. Each tool call is executed through the OpenAI function-calling interface; the language model emits a JSON stub, the back-end runs the call, and the retrieved snippets flow into a transient database before being re-injected into the prompt. Finally, a frozen GPT-4o model synthesizes those multi-modal fragments with the original question, streams an Augmented Response to the UI, and stores the turn in memory, readying the loop for whatever the user asks next. Responses are fast, as the tool calling system is a series of quick API calls.
@@ 4. Data Modes And Tools
For every security in the user’s portfolio, we ingest MSCI MAC(shepard2020msci) factor scores and the accompanying methodology text. Before embedding, each record is wrapped in a context shell (see section 5): a short paragraph that “tokenizes” the numerical z‑scores within natural‑language sentences so the LLM can reason about their meaning rather than seeing them as opaque numbers. The resulting embeddings power the get_factors tool.
@@
The final mode is a multi‑entity knowledge graph that links Company, Product, Input Product, Input to Input Product, Industry, and Location nodes via relations such as Produces, Has Input, Manufactured In and so forth (the full list of edge types are given in LABEL:tab:entity_rships_kg in A. Although illustrative rather than fully industrial, the graph is constructed with a pipeline akin to AIPNet (fetzer2024aipnet), to query ChatGPT’s intrinsic knowledge base to create a dataset of the desired form. These paths enter the prompt as restructured text within a context shell, exposing the semantic relationships between nodes on the traversed graph.
@@ 5. Context Shell
Domain risk analysis is based on numbers: z-scored factor exposures, portfolio weights, valuation ratios. Large language models, however, treat raw figures as opaque tokens and rarely attach contextual economic meaning to them. To retain a frozen, generic LLM while still letting it reason quantitatively, our proposed system wraps every table row from the MSCI Multi-Asset-Class (MAC) factor model in a context shell: a short paragraph that “tokenizes” each figure inside an explanatory sentence so that neighboring words endow the number with semantics. The surrounding text becomes part of the embedding; the number itself is now a first-class token that via attention relationships with pertinent context becomes easily interpretable. A related but different idea has been explored in (anthropic2024contextual).
@@ 6. Network‑Science Path Discovery
In practice supply chain networks follow a scale-free, power-law degree distribution (mungo2023reconstruct): hubs such as Electronic Integrated Circuits in the middle of Figure 3 in the Apple network expose a wealth of information in just one hop, whereas peripheral nodes require additional traversal hops for useful coverage.